### v1

@@ 5.1 Fibrations with base $\mathbb{P}^{2}$
In §4.3.1, we enumerated fibrations over the sixteen weak Fano bases. Note that one of the Fano bases, MT ID 1, is simply $\mathbb{P}^{2}$. Automorphism-inequivalent toric fibrations over $\mathbb{P}^{2}$ were enumerated in Braun:2011ux; our count of 102,581 such fibrations exactly matches with the count in that work. [^9] We thank V. Braun for helpful discussions on this comparison. In the following subsections, and in particular 5.3, we will re-analyze some specific examples discussed in Braun:2011ux and provide a new perspective on their physics and geometry.
@@ 5.2.1 SU(2) on non-toric divisors
monomials, forming a generic homogeneous polynomial of degree $n$ on $\mathbb{P}^{2}$. (As discussed in § 2.6, monomials in $\beta$ correspond to points (2,1;m)(2,-1;m) in the dual polytope, which are easily enumerated.) Thus the fibrations of these polytopes contain an $SU(2)$ gauge group that has been tuned over a degree-$n$ curve in the base, or equivalently on the nontoric divisor $nH$. From anomaly cancellation (see e.g. Table 7 in Johnson:2016qar), we know that the massless matter spectrum charged under such an SU(2) consists of $6n^{2}+16(1-g)$ fields in the fundamental representation and $g$ fields in the adjoint representation, where g=(n1)(n2)/2g=(n-1)(n-2)/2 is the genus of the generic degree-$n$ curve. Recalling that the Cartan elements of the adjoint do not contribute to $H_{\text{charged}}$, we see that this matter spectrum nicely matches with (14c) and the above formula for $h^{2,1}$.
@@
Next, we consider the case $n=9$, for which Eq. 39 once again furnishes a reflexive polytope, with Hodge numbers $h^{1,1}(X)=4$, $h^{2,1}(X)=112$. As before, this polytope has a single fibration, a nonstandard stacking of $\mathbb{P}^{2}$ over of $F_{10}$, and for $n=9$, we find that $\beta$ consists of 55 monomials, as expected for $\mathfrak{su}$(2) hosted on a degree-nine curve. From the above discussion, for $\mathfrak{su}$(2) on a nonic curve we expect 28 adjoints and 54 fundamentals, which would yield $h^{2,1}(X)=111$. To satisfy Eq. 13, we require an additional gauge factor; since no toric divisors host nonabelian gauge groups and $\alpha$ consists of a single monomial, we conclude that the missing gauge factor is a $U(1)$. Furthermore, since the discrepancies in the Hodge numbers in the model without the U(1) factor are the same, all matter charged under this abelian gauge factor must also be charged under the SU(2). These constraints are nicely satisfied by the theory with gauge group $(SU(2)\times U(1))/\mathbb{Z}_{2}$ , in which there are 27 fields in each of the representations $({\bf 2},\pm{\bf 1})$. This theory illustrates two general principles that have been identified in 6D F-theory models, and are hypothesized to hold in general but have not yet been fully proven. First, the global structure of the gauge group is fixed by the massless charge completeness hypothesis Morrison:2021wuv, which suggests that the gauge group does not contain any central elements under which the massless spectrum is invariant; this spectrum is invariant under the $\mathbb{Z}_{2}$ diagonal central element of $SU(2)\times U(1)$, so we select the quotient to project out this element. Second, the automatic enhancement conjecture Raghuram:2020vxm states that the additional U(1) factor here is a necessary consequence of the SU(2) spectrum; that is, since the U(1) factor cannot be broken through a Higgs mechanism while preserving the SU(2) symmetry, any Weierstrass model over $\mathbb{P}^{2}$ with an SU(2) tuned over a degree 9 curve will automatically force the presence of an additional U(1) factor of this form that enhances the Mordell-Weil group of the associated threefold. Indeed, precisely this example (as well as the cases with $n=10,11,12$) is discussed in Raghuram:2020vxm.
@@
It is also worth noting that for the degree 12 case, the spectrum consists of 55 adjoints, and the gauge group is actually $SO(3)=SU(2)/\mathbb{Z}_{2}$. This matches the massless charge completeness hypothesis and is shown explicitly by constructing the explicit torsion section in this case in Morrison:2021wuv.
@@ 5.2.2 SU(4) on a non-toric divisor
This polytope has $h^{1,1}=6,$ $h^{2,1}=60$, and is a stacking of $\mathbb{P}^{2}$ over the vertex $(1,0)$ of the fiber $F_{13}$. The analysis of Klevers:2014bqa indicates that $SU(4)$ is tuned on the vanishing locus of the coefficient $a_{012}$ in the cubic; in this geometry, $a_{012}$ contains 28 monomials, as expected for a generic polynomial of degree 6, so we conclude that this fibration has an $SU(4)$ gauge group on a sextic in $\mathbb{P}^{2}$. From anomaly cancellation, we can ascertain that the massless spectrum for such an SU(4) factor will consist of no fields in the fundamental, 18 fields in the antisymmetric (6) representation, and 10 fields in the adjoint (15) representation. If this was the only gauge group and matter, we would expect to have a Calabi-Yau threefold with $h^{1,1}=5,h^{2,1}=59$. Thus, there is an additional hidden gauge factor. Since there are no other nonabelian factors manifest in the Weierstrass model, this must again be an abelian factor. Furthermore, since the discrepancies in the Hodge numbers are the same, again, all matter charged under this abelian gauge factor must also be charged under the nonabelian factor SU(4). These constraints are nicely satisfied by the theory with gauge group
@@ 5.2.3 SU(3) on a non-toric curve
containing 18 fields in each of the representations $({\bf 3},{\bf 1},{\bf 1}),({\bf 3},{\bf-1},{\bf 0}),({\bf 3},{\bf 0},{\bf-1})$. This spectrum again manifests the principles of massless charge completeness and automatic enhancement; it seems that any tuning of SU(3) on a degree 6 curve in $\mathbb{P}^{2}$ will automatically have these extra U(1) factors in this form. Note that this spectrum can also be understood as arising from a Higgsing of a theory with gauge group $(SU(3)\times SU(3))/\mathbb{Z}_{3}$, with the SU(3) factors respectively living on a sextic and a cubic, where Higgsing is carried out on the single adjoint representation of the second SU(3) factor. This construction of the above $U(1)\times U(1)$ spectrum was realized in Cvetic:2015ioa.
@@ 5.3 SCFTs and non-toric bases; Example: Maximum rank increase over $\mathbb{P}^{2}$
It is interesting to consider the example of the fibration over $\mathbb{P}^{2}$ with the largest $h^{1,1}(X)$. This Calabi Yau was identified in Braun:2011ux, where it was noted that the apparent gauge group is $SU(27)$, supported on the divisor $H$, and there are an additional 84 divisors encoded in non-flat fibers, giving a total of $h^{1,1}(X)=112$. The polytope, whose ID is v05-65, has $h^{2,1}=4$, and the fiber type in this fibration is $F_{16}$. We can understand this geometry in terms of an interesting SCFT with a natural geometric resolution.
Simply from the structure of the gauge group and an analysis of the F-theory and Weierstrass models, it is known that the gauge group SU(27) cannot be realized on the base $\mathbb{P}^{2}$ in a straightforward way. Anomaly cancellation considerations show that with only the gauge group SU($N$), for $N>24$ the number of fundamental matter representations would go negative, and SU(24) is the largest SU($N$) group that can be tuned in an F-theory Weierstrass model without singularities outside the Kodaira classification of Table 1 Morrison:2011mb.
Indeed, as pointed out in Braun:2011ux, in this model, there are three points on the base where the orders of vanishing of the Weierstrass coefficient $f,g$ reach (4, 6), signaling a non-flat fiber, which can be interpreted as an SCFT coupled to the gravity theory HeckmanMorrisonVafa; Heckman:2015bfa.
@@
In the final resolved geometry, we have a non-toric (but semi-toric MartiniTaylorSemitoric) base with $h^{1,1}(B)=10$. The gauge group is SU(27) $\times$ (SU(18) $\times$ SU(9))3, and all seven SU($N$) factors live on curves of self-intersection -2. In general, the matter content for an SU($N$) gauge factor on a -2 curve is $2N$ fundamental matter fields. At each intersection between $\mathrm{SU}(N)\times\mathrm{SU}(N^{\prime})$ a bifundamental matter field is supported. It is easy to check that these bifundamentals exactly saturate the anomaly cancellation condition for each $\mathrm{SU}(N)$ factor. The total number of charged matter fields is $3\times 18\times(9+27)=1944.$ We have $T=9$, and there are $V=1937$ vector multiplets, from which it is easy to confirm that all anomaly cancellation conditions are properly satisfied with $h^{2,1}=4$.
This example illustrates a general feature of many fibrations in the database: in the elliptically fibered phase, there are non-flat fibers associated with (4, 6) fibration singularities in the Weierstrass model, corresponding in the F-theory picture to local strongly coupled superconformal field theory sectors. There is a resolved (non-toric) phase for such Calabi-Yau threefolds in which points on the base are blown up, giving generically a non-toric base, often with an enhanced gauge group.  [^11] Note that similar things can happen for different flop phases/triangulations of a CY where both bases are toric; for example, for a specific triangulation of the generic elliptic fibration (simple stacking) over Hirzebruch $\mathbb{F}_{1}$, the resulting CY3 is elliptically fibered over $\mathbb{P}^{2}$ with a non-flat fiber over the toric point that would be blown up to form $\mathbb{F}_{1}$ rw-forthcoming; we expect that many examples of this that connect to a non-toric base phase similarly can be associated with flop transitions in the extended Kähler cone. This is a much more dramatic version of the familiar presence of (4, 6) loci on -9, -10, and -11 curves in a toric base. The structures of this kind that arise in the toric hypersurface database seem to describe elliptic fibrations over a wide range of non-toric bases. We have not explored this in any depth, but leave for further study the range of non-toric bases that are realized in this fashion in the database. It would be interesting to compare the set of bases realized in this way with previous partial analyses of the full set of non-toric bases MartiniTaylorSemitoric; TaylorWangNon-toric.
@@ 5.4 The most common base (base 72) and fiber ($F_{6}$)
Thus, whenever $a_{102}$ consists of more than one monomial, we have a Tate-tuned $\mathfrak{su}(2)$ over the locus $a_{102}=0$ Klevers:2014bqa. Additionally, it was shown in Klevers:2014bqa that $F_{6}$ fibrations always have at least a $U(1)$ abelian sector; whereas the “generic” $\mathfrak{su}(2)$ might or might not be realized in any particular example, this $U(1)$ always appears, regardless of whether any of the $a_{ijk}$ consist of a single monomial. The presence of this extra Mordell-Weil section can also be seen immediately from the fact that the fiber $F_{6}$ contains two distinct -1 curves, as mentioned in Section 2.5
@@ 5.4.1 A Typical (72,6) Fibration
Because base 72 has $h^{1,1}(B)=5$, by STW we have a gauge group of total rank 2051=1420-5-1=14. Moreover, by the analysis of Klevers:2014bqa, the abelian sector has rank at least one, so the total rank of the nonabelian sector of the gauge group is no larger than 13.
@@ 5.4.2 The Largest $h^{2,1}$ (72,6) Fibration
By Shioda-Tate-Wazir, we see that the total rank of the gauge group is $r=h^{1,1}(X)-h^{1,1}(B)-1=1.$ On the other hand, all $F_{6}$ fibrations have at least one $U(1)$ factor, so in this example the entire gauge group is the universal $U(1)$ identified in Klevers:2014bqa, with no nonabelian sector. Indeed, the Weierstrass coefficient $a_{102}$ which gives rise to the universal $SU(2)$ factor consists of only a single monomial, and hence the $SU(2)$ factor does not appear; moreover, none of the $a_{ijk}$ vanish on any toric divisor, and hence we have no toric Tate tunings, as expected.
@@ 5.4.3 The Largest $h^{1,1}$ (72,6) Fibration
where $z=0$ is a local description of the toric divisor and $n,m>0$ depend on the particular divisor. Thus, if the divisor $D$ intersects the zero locus of $a_{111}$, we then have a $(4,6)$ point, which must be blown up. $a_{111}$ is a section of K-K Klevers:2014bqa, so it does not intersect the -2 curves, each of which hosts an $\mathfrak{su}(10)$ gauge algebra. Instead, we have $(4,6)$ points on the divisors (1,2),(1,0)(-1,-2),(1,0) and (1,0)(-1,0), each of which must be blown up.
@@
We now pass to anomaly cancellation, assuming that the gauge group is as in 120b. We have
@@ 5.5 Fiber $F_{11}$ and the Standard Model
As indicated in Table 2, the generic gauge group for fibrations with fiber $F_{11}$ is $(U(1)\times SU(2)\times SU(3))/\mathbb{Z}_{6}$, i.e. the Standard Model gauge group; this observation has generated significant interest in this fiber Klevers:2014bqa, and is the origin of, e.g., the quadrillion Standard Models of Cvetic:2019gnh. This arises because $F_{11}$ can be realized by adding a vertex to $F_{10}$ so that its defining polynomial is a specialization of Eq. 9, namely
@@
While we obviously do not live in six space-time dimensions, we expect that any special feature of this fiber type would likely be shared between Calabi-Yau threefolds and the elliptic Calabi-Yau fourfolds that are used for compactifications to 4D, and so it is natural to wonder whether this structure is in some way favored or special in the F-theory landscape.  [^13] We would like to thank Damien Mayorga Pena for discussions on this issue. Like the other fibers, $F_{11}$ captures a particular slice of the moduli space of Weierstrass models, with certain specific tuned structures. The general form of the Weierstrass model with this tuned gauge group was described in Taylor:2019wnm; Raghuram:2019efb. From this point of view, there is no a priori reason to expect this class of fibrations to be particularly favored, but it is certainly a good question to investigate.
@@
The universal gauge group contribution is not the only way to realize an exact SM in $F_{11}$ fibrations. Even if the nontoric divisors $\alpha,\beta$ do not host gauge factors in a given fibration, the missing gauge factors can be realized on toric divisors; in such examples, the $F_{11}$ can effectively be thought of as providing the $U(1)$ factor, which is what makes it more useful than $F_{10}$ in the context of the SM. Note, however that in this context the global structure of the gauge group will generally not include the $\mathbb{Z}_{6}$ quotient, which in the 4D context likely compromises any natural realization of Standard Model-like matter in such constructions Taylor:2019ots.

### v2

@@ 5.1 Fibrations with base $\mathbb{P}^{2}$
In §4.3.1, we enumerated fibrations over the sixteen weak Fano bases. Note that one of the Fano bases, MT ID 1, is simply $\mathbb{P}^{2}$. Automorphism-inequivalent toric fibrations over $\mathbb{P}^{2}$ were enumerated in Braun (2013); our count of 102,581 such fibrations exactly matches with the count in that work. [^11] We thank V. Braun for helpful discussions on this comparison. In the following subsections, and in particular 5.3, we will re-analyze some specific examples discussed in Braun (2013) and provide a new perspective on their physics and geometry.
@@ 5.2.1 SU(2) on non-toric divisors
monomials, forming a generic homogeneous polynomial of degree $n$ on $\mathbb{P}^{2}$. (As discussed in § 2.6, monomials in $\beta$ correspond to points (2,1;m)(2,-1;m) in the dual polytope, which are easily enumerated.) Thus the fibrations of these polytopes contain an $SU(2)$ gauge group that has been tuned over a degree-$n$ curve in the base, or equivalently on the nontoric divisor $nH$. From anomaly cancellation (see e.g. Table 7 in Johnson and Taylor (2016)), we know that the massless matter spectrum charged under such an SU(2) consists of $6n^{2}+16(1-g)$ fields in the fundamental representation and $g$ fields in the adjoint representation, where g=(n1)(n2)/2g=(n-1)(n-2)/2 is the genus of the generic degree-$n$ curve. Recalling that the Cartan elements of the adjoint do not contribute to $H_{\text{charged}}$, we see that this matter spectrum nicely matches with (15c) and the above formula for $h^{2,1}$.
@@
Next, we consider the case $n=9$, for which Eq. 40 once again furnishes a reflexive polytope, with Hodge numbers $h^{1,1}(X)=4$, $h^{2,1}(X)=112$. As before, this polytope has a single fibration, a nonstandard stacking of $\mathbb{P}^{2}$ over of $F_{10}$, and for $n=9$, we find that $\beta$ consists of 55 monomials, as expected for $\mathfrak{su}$(2) hosted on a degree-nine curve. From the above discussion, for $\mathfrak{su}$(2) on a nonic curve we expect 28 adjoints and 54 fundamentals, which would yield $h^{2,1}(X)=111$. To satisfy Eq. 14, we require an additional gauge factor; since no toric divisors host nonabelian gauge groups and $\alpha$ consists of a single monomial, we conclude that the missing gauge factor is a $U(1)$. Furthermore, since the discrepancies in the Hodge numbers in the model without the U(1) factor are the same, all matter charged under this abelian gauge factor must also be charged under the SU(2). These constraints are nicely satisfied by the theory with gauge group $(SU(2)\times U(1))/\mathbb{Z}_{2}$ , in which there are 27 fields in each of the representations $({\bf 2},\pm{\bf 1})$. This theory illustrates two general principles that have been identified in 6D F-theory models, and are hypothesized to hold in general but have not yet been fully proven. First, the global structure of the gauge group is fixed by the massless charge completeness hypothesis Morrison and Taylor (2021), which suggests that the gauge group does not contain any central elements under which the massless spectrum is invariant; this spectrum is invariant under the $\mathbb{Z}_{2}$ diagonal central element of $SU(2)\times U(1)$, so we select the quotient to project out this element. Second, the automatic enhancement conjecture Raghuram et al. (2021) states that the additional U(1) factor here is a necessary consequence of the SU(2) spectrum; that is, since the U(1) factor cannot be broken through a Higgs mechanism while preserving the SU(2) symmetry, any Weierstrass model over $\mathbb{P}^{2}$ with an SU(2) tuned over a degree 9 curve will automatically force the presence of an additional U(1) factor of this form that enhances the Mordell-Weil group of the associated threefold. Indeed, precisely this example (as well as the cases with $n=10,11,12$) is discussed in Raghuram et al. (2021).
@@
It is also worth noting that for the degree 12 case, the spectrum consists of 55 adjoints, and the gauge group is actually $SO(3)=SU(2)/\mathbb{Z}_{2}$. This matches the massless charge completeness hypothesis and is shown explicitly by constructing the explicit torsion section in this case in Morrison and Taylor (2021).
@@ 5.2.2 SU(4) on a non-toric divisor
This polytope has $h^{1,1}=6,$ $h^{2,1}=60$, and is a stacking of $\mathbb{P}^{2}$ over the vertex $(1,0)$ of the fiber $F_{13}$. The analysis of Klevers et al. (2015) indicates that $SU(4)$ is tuned on the vanishing locus of the coefficient $a_{012}$ in the cubic; in this geometry, $a_{012}$ contains 28 monomials, as expected for a generic polynomial of degree 6, so we conclude that this fibration has an $SU(4)$ gauge group on a sextic in $\mathbb{P}^{2}$. From anomaly cancellation, we can ascertain that the massless spectrum for such an SU(4) factor will consist of no fields in the fundamental, 18 fields in the antisymmetric (6) representation, and 10 fields in the adjoint (15) representation. If this was the only gauge group and matter, we would expect to have a Calabi-Yau threefold with $h^{1,1}=5,h^{2,1}=59$. Thus, there is an additional hidden gauge factor. Since there are no other nonabelian factors manifest in the Weierstrass model, this must again be an abelian factor. Furthermore, since the discrepancies in the Hodge numbers are the same, again, all matter charged under this abelian gauge factor must also be charged under the nonabelian factor SU(4). These constraints are nicely satisfied by the theory with gauge group
@@ 5.2.3 SU(3) on a non-toric curve
containing 18 fields in each of the representations $({\bf 3},{\bf 1},{\bf 1}),({\bf 3},{\bf-1},{\bf 0}),({\bf 3},{\bf 0},{\bf-1})$. This spectrum again manifests the principles of massless charge completeness and automatic enhancement; it seems that any tuning of SU(3) on a degree 6 curve in $\mathbb{P}^{2}$ will automatically have these extra U(1) factors in this form. Note that this spectrum can also be understood as arising from a Higgsing of a theory with gauge group $(SU(3)\times SU(3))/\mathbb{Z}_{3}$, with the SU(3) factors respectively living on a sextic and a cubic, where Higgsing is carried out on the single adjoint representation of the second SU(3) factor. This construction of the above $U(1)\times U(1)$ spectrum was realized in Cvetič et al. (2015).
@@ 5.3 SCFTs and non-toric bases; Example: Maximum rank increase over $\mathbb{P}^{2}$
It is interesting to consider the example of the fibration over $\mathbb{P}^{2}$ with the largest $h^{1,1}(X)$. This Calabi Yau was identified in Braun (2013), where it was noted that the apparent gauge group is $SU(27)$, supported on the divisor $H$, and there are an additional 84 divisors encoded in non-flat fibers, giving a total of $h^{1,1}(X)=112$. The polytope, whose ID is v05-65, has $h^{2,1}=4$, and the fiber type in this fibration is $F_{16}$. We can understand this geometry in terms of an interesting SCFT with a natural geometric resolution.
Simply from the structure of the gauge group and an analysis of the F-theory and Weierstrass models, it is known that the gauge group SU(27) cannot be realized on the base $\mathbb{P}^{2}$ in a straightforward way. Anomaly cancellation considerations show that with only the gauge group SU($N$), for $N>24$ the number of fundamental matter representations would go negative, and SU(24) is the largest SU($N$) group that can be tuned in an F-theory Weierstrass model without singularities outside the Kodaira classification of Table 1 Morrison and Taylor (2012b).
Indeed, as pointed out in Braun (2013), in this model, there are three points on the base where the orders of vanishing of the Weierstrass coefficient $f,g$ reach (4, 6), signaling a non-flat fiber, which can be interpreted as an SCFT coupled to the gravity theory Heckman et al. (2014, 2015).
@@
In the final resolved geometry, we have a non-toric (but semi-toric Martini and Taylor (2015)) base with $h^{1,1}(B)=10$. The gauge group is SU(27) $\times$ (SU(18) $\times$ SU(9))3, and all seven SU($N$) factors live on curves of self-intersection -2. In general, the matter content for an SU($N$) gauge factor on a -2 curve is $2N$ fundamental matter fields. At each intersection between $\mathrm{SU}(N)\times\mathrm{SU}(N^{\prime})$ a bifundamental matter field is supported. It is easy to check that these bifundamentals exactly saturate the anomaly cancellation condition for each $\mathrm{SU}(N)$ factor. The total number of charged matter fields is $3\times 18\times(9+27)=1944.$ We have $T=9$, and there are $V=1937$ vector multiplets, from which it is easy to confirm that all anomaly cancellation conditions are properly satisfied with $h^{2,1}=4$.
This example illustrates a general feature of many fibrations in the database: in the elliptically fibered phase, there are non-flat fibers associated with (4, 6) fibration singularities in the Weierstrass model, corresponding in the F-theory picture to local strongly coupled superconformal field theory sectors. There is a resolved (non-toric) phase for such Calabi-Yau threefolds in which points on the base are blown up, giving generically a non-toric base, often with an enhanced gauge group.  [^13] Note that similar things can happen for different flop phases/triangulations of a CY where both bases are toric; for example, for a specific triangulation of the generic elliptic fibration (simple stacking) over Hirzebruch $\mathbb{F}_{1}$, the resulting CY3 is elliptically fibered over $\mathbb{P}^{2}$ with a non-flat fiber over the toric point that would be blown up to form $\mathbb{F}_{1}$ Nally and Taylor ; we expect that many examples of this that connect to a non-toric base phase similarly can be associated with flop transitions in the extended Kähler cone. This is a much more dramatic version of the familiar presence of (4, 6) loci on -9, -10, and -11 curves in a toric base. The structures of this kind that arise in the toric hypersurface database seem to describe elliptic fibrations over a wide range of non-toric bases. We have not explored this in any depth, but leave for further study the range of non-toric bases that are realized in this fashion in the database. It would be interesting to compare the set of bases realized in this way with previous partial analyses of the full set of non-toric bases Martini and Taylor (2015); Taylor and Wang (2017).
@@ 5.4 The most common base (base 72) and fiber ($F_{6}$)
Thus, whenever $a_{102}$ consists of more than one monomial, we have a Tate-tuned $\mathfrak{su}(2)$ over the locus $a_{102}=0$ Klevers et al. (2015). Additionally, it was shown in Klevers et al. (2015) that $F_{6}$ fibrations always have at least a $U(1)$ abelian sector; whereas the “generic” $\mathfrak{su}(2)$ might or might not be realized in any particular example, this $U(1)$ always appears, regardless of whether any of the $a_{ijk}$ consist of a single monomial. The presence of this extra Mordell-Weil section can also be seen immediately from the fact that the fiber $F_{6}$ contains two distinct -1 curves, as mentioned in Section 2.5
@@ 5.4.1 A Typical (72,6) Fibration
Because base 72 has $h^{1,1}(B)=5$, by STW we have a gauge group of total rank 2051=1420-5-1=14. Moreover, by the analysis of Klevers et al. (2015), the abelian sector has rank at least one, so the total rank of the nonabelian sector of the gauge group is no larger than 13.
@@ 5.4.2 The Largest $h^{2,1}$ (72,6) Fibration
By Shioda-Tate-Wazir, we see that the total rank of the gauge group is $r=h^{1,1}(X)-h^{1,1}(B)-1=1.$ On the other hand, all $F_{6}$ fibrations have at least one $U(1)$ factor, so in this example the entire gauge group is the universal $U(1)$ identified in Klevers et al. (2015), with no nonabelian sector. Indeed, the Weierstrass coefficient $a_{102}$ which gives rise to the universal $SU(2)$ factor consists of only a single monomial, and hence the $SU(2)$ factor does not appear; moreover, none of the $a_{ijk}$ vanish on any toric divisor, and hence we have no toric Tate tunings, as expected.
@@ 5.4.3 The Largest $h^{1,1}$ (72,6) Fibration
where $z=0$ is a local description of the toric divisor and $n,m>0$ depend on the particular divisor. Thus, if the divisor $D$ intersects the zero locus of $a_{111}$, we then have a $(4,6)$ point, which must be blown up. $a_{111}$ is a section of K-K Klevers et al. (2015), so it does not intersect the -2 curves, each of which hosts an $\mathfrak{su}(10)$ gauge algebra. Instead, we have $(4,6)$ points on the divisors (1,2),(1,0)(-1,-2),(1,0) and (1,0)(-1,0), each of which must be blown up.
@@
We now pass to anomaly cancellation, assuming that the gauge group is as in 121b. We have
@@ 5.5 Fiber $F_{11}$ and the Standard Model
As indicated in Table 2, the generic gauge group for fibrations with fiber $F_{11}$ is $(U(1)\times SU(2)\times SU(3))/\mathbb{Z}_{6}$, i.e. the Standard Model gauge group; this observation has generated significant interest in this fiber Klevers et al. (2015), and is the origin of, e.g., the quadrillion Standard Models of Cvetič et al. (2019). This arises because $F_{11}$ can be realized by adding a vertex to $F_{10}$ so that its defining polynomial is a specialization of Eq. 10, namely
@@
While we obviously do not live in six space-time dimensions, we expect that any special feature of this fiber type would likely be shared between Calabi-Yau threefolds and the elliptic Calabi-Yau fourfolds that are used for compactifications to 4D, and so it is natural to wonder whether this structure is in some way favored or special in the F-theory landscape.  [^15] We would like to thank Damien Mayorga Pena for discussions on this issue. Like the other fibers, $F_{11}$ captures a particular slice of the moduli space of Weierstrass models, with certain specific tuned structures. The general form of the Weierstrass model with this tuned gauge group was described in Taylor and Turner (2020); Raghuram et al. (2020). From this point of view, there is no a priori reason to expect this class of fibrations to be particularly favored, but it is certainly a good question to investigate.
@@
The universal gauge group contribution is not the only way to realize an exact SM in $F_{11}$ fibrations. Even if the nontoric divisors $\alpha,\beta$ do not host gauge factors in a given fibration, the missing gauge factors can be realized on toric divisors; in such examples, the $F_{11}$ can effectively be thought of as providing the $U(1)$ factor, which is what makes it more useful than $F_{10}$ in the context of the SM. Note, however that in this context the global structure of the gauge group will generally not include the $\mathbb{Z}_{6}$ quotient, which in the 4D context likely compromises any natural realization of Standard Model-like matter in such constructions Taylor and Turner (2019).